OpenStudy (anonymous):

One number is five more than another, and their sum is three less than three times the smaller. Find the numbers. If x represents the smaller number, which equation could be used to solve for x? x 2 + 5 = 3x - 3 2x + 5 = 3x - 3 2x + 5 = 3(x - 3)

OpenStudy (anonymous):

Let x = smaller number Let x + 5 = bigger number smaller number PLUS bigger number = 3 times smaller number minus three. Here's what the statement looks like as an equation: x + x + 5 = 3x - 3 Now, solve for x. 2x + 5 = 3x - 3 2x - 3x = -3 -5 -x = - 8 x = 8...smaller number x + 5 becomes 8 + 5 = 13...bigger number The smaller number is 8 and and bigger number is 13.